# VL-LN Bench: Towards Long-horizon Goal-oriented Navigation with Active Dialogs

### v1

@@ IV-B Result
Why Instance Object Navigation is Hard. Tab. II shows that, even when trained on instance-goal data, performance remains far below prior goal-oriented benchmarks (e.g., VLLN-O: 59.3% SR on ObjectNav). We attribute this gap to two factors. First, long-horizon exploration: in ObjectNav, reaching any instance of a category (e.g., any of the seven “Chair” candidates in Fig. 1) suffices, whereas instance-goal navigation requires locating a specific instance, often necessitating substantially longer exploration. Second, agents often struggle to detect the target instance because attribute–image alignment is challenging; they may overlook or misalign attributes, leading to stops at same-category distractors on the true target.
@@
Dialog Helps. Across both IION and ION, VLLN-D achieves the best results, demonstrating the value of proactive querying while leaving room for further improvement. Dialog in the ambiguity setting (IION) is more effective than in the disambiguation setting (ION) (MSP: 2.76 vs. 2.16). To better analyze the role of dialog during instance object navigation, we categorize failures into three types: Exploration Fail, Detection Fail, and Stop Fail. Exploration Fail occurs when the agent keeps acting until the maximum number of steps without ever entering the oracle-success radius ($OS=0$). Unlike prior goal-oriented navigation settings, we further split Detection Fail into Wrong Detection and Ambiguity: the former happens when the agent reaches the correct region ($OS=1$) but fails to recognize the target instance, while the latter occurs when it stops at an instance of the same category as the target. Stop Fail denotes cases where the agent stops near the target (<1m<1m) but remains beyond the stop threshold (>0.25m>0.25m). As shown in Tab. III, when the agent is allowed to query an oracle, exploration failures decrease substantially: from 89 to 71 in IION and from 84 to 46 in ION. This demonstrates that the agent can leverage information obtained through dialog to improve exploration efficiency. Notably, as shown in the supplementary video, the agent is able to ground natural-language short-route guidance into executable actions. We attribute it to co-training with VLN data, which, although different from goal-oriented navigation, teaches the model to follow natural-language route descriptions.
@@
Moreover, dialog helps resolve ambiguity. In both IION and ION settings, ambiguity-related failures decrease once the dialog is enabled. In ION, the reduction from 143 to 124 primarily stems from disambiguation questions that guide the agent away from incorrect instances, as the attributes necessary to verify the target instance are already available, rendering additional attribute-related queries less informative. For IION, both attribute and disambiguation questions contribute to eliminating task ambiguity, make the ambiguity failures drop from 143 to 124.
@@
Key Challenges and Directions. To better understand IION’s challenges, we additionally evaluate four interaction settings on 100 episodes randomly sampled from the test set: Human-Human, Human-Oracle, Agent-Human, and Agent-Oracle. Results are reported in Tab. IV. First, we argue that image–attribute alignment is the primary bottleneck in both the ION and IION tasks. Even with full instructions and dialog capability, the success rate remains 25% (shown in Tab. II), and 73% of failures (shown in Tab. III) are due to detection errors. A promising direction is to train with hard negatives such as same-category objects with different attributes, which promotes instance-level discrimination and stronger attribute grounding. Zero-shot methods face the same challenge. Their performance hinges on detectors that can align detailed instance descriptions with visual evidence.
@@
Secondly, the agent’s questioning ability remains limited. Reliably disambiguating a target instance from same-category distractors through dialog is still difficult. VLLN-D performs worse on IION (20.2%) than VLLN-I on ION (22.4%), indicating that dialog offers less guidance than a full instruction. We also find that a gap persists between agent-driven and human-driven proactive interaction. Human-Human reaches 93% SR with only about two questions on average, which suggests that a small number of well-chosen queries is sufficient to complete the task. Achieving such behavior requires the agent to exploit its observation history to select maximally informative questions that either guide exploration or shrink the candidate set, placing stronger demands on grounding, planning, and reasoning.
@@ IV-C Additional Insights
Reliability of evaluation and Oracle The benchmark provides a reliable testbed for IION, as humans achieve high success rates. The oracle is also dependable, with the Human–Oracle setting attaining performance comparable to Human–Human (91% vs. 93%), though at the cost of more dialog turns (9.72 vs. 2.04 on average). This discrepancy arises for two reasons: some queries fall outside the oracle’s knowledge or response schema, and users interacting with AI systems often seek additional confirmation (e.g., asking a disambiguation question even after the target has been identified). Although not flawless, the oracle is sufficient for evaluating IION, as performance in the Agent–Human and Agent–Oracle settings is similar (16% vs. 17%).
@@ V Conclusion
This paper investigates the Interactive Instance Object Navigation (IION) task and introduces the VL-LN benchmark, which includes a long-horizon dataset comprising $\sim$ 41k automatically collected dialogue-augmented trajectories for training, along with an evaluation protocol involving an oracle for agent assessment. This benchmark enables agents to explore long-horizon environments and engage in meaningful dialogs. Our experiments demonstrate that incorporating active dialog significantly improves performance in both the IION and ION tasks, achieving state-of-the-art results. Additionally, we highlight key challenges in the IION task and provide some interesting findings from our experiments.

### v3

@@ IV-B Result
Why Instance Goal Navigation is Hard. Tab. II shows that, even when trained on instance-goal data, performance remains far below prior goal-oriented benchmarks (e.g., VLLN-O: 59.3% SR on ObjectNav). We attribute this gap to two factors. First, long-horizon exploration: in ObjectNav, reaching any instance of a category (e.g., any of the seven “Chair” candidates in Fig. 1) suffices, whereas instance-goal navigation requires locating a specific instance, often necessitating substantially longer exploration. Second, agents often struggle to detect the target instance because attribute–image alignment is challenging; they may overlook or misalign attributes, leading to stops at same-category distractors on the true target.
@@
Dialog Helps. Across both IIGN and IGN, VLLN-D achieves the best results, demonstrating the value of proactive querying while leaving room for further improvement. Dialog in the ambiguity setting (IIGN) is more effective than in the disambiguation setting (IGN) (MSP: 2.76 vs. 2.16). To better analyze the role of dialog during instance goal navigation, we categorize failures into three types: Exploration Fail, Detection Fail, and Stop Fail. Exploration Fail occurs when the agent keeps acting until the maximum number of steps without ever entering the oracle-success radius ($OS=0$). Unlike prior goal-oriented navigation settings, we further split Detection Fail into Wrong Detection and Ambiguity: the former happens when the agent reaches the correct region ($OS=1$) but fails to recognize the target instance, while the latter occurs when it stops at an instance of the same category as the target. Stop Fail denotes cases where the agent stops near the target (<1m<1m) but remains beyond the stop threshold (>0.25m>0.25m). As shown in Tab. III, when the agent is allowed to query an oracle, exploration failures decrease substantially: from 89 to 71 in IIGN and from 84 to 46 in IGN. This demonstrates that the agent can leverage information obtained through dialog to improve exploration efficiency. Notably, as shown in the supplementary video, the agent is able to ground natural-language short-route guidance into executable actions. We attribute it to co-training with VLN data, which, although different from goal-oriented navigation, teaches the model to follow natural-language route descriptions.
@@
Moreover, dialog helps resolve ambiguity. In both IIGN and IGN settings, ambiguity-related failures decrease once the dialog is enabled. In IGN, the reduction from 143 to 124 primarily stems from disambiguation questions that guide the agent away from incorrect instances, as the attributes necessary to verify the target instance are already available, rendering additional attribute-related queries less informative. For IIGN, both attribute and disambiguation questions contribute to eliminating task ambiguity, make the ambiguity failures drop from 143 to 124.
@@
Key Challenges and Directions. To better understand IIGN’s challenges, we additionally evaluate four interaction settings on 100 episodes randomly sampled from the test set: Human-Human, Human-Oracle, Agent-Human, and Agent-Oracle. Results are reported in Tab. IV. First, we argue that image–attribute alignment is the primary bottleneck in both the IGN and IIGN tasks. Even with full instructions and dialog capability, the success rate remains 25% (shown in Tab. II), and 73% of failures (shown in Tab. III) are due to detection errors. A promising direction is to train with hard negatives such as same-category objects with different attributes, which promotes instance-level discrimination and stronger attribute grounding. Zero-shot methods face the same challenge. Their performance hinges on detectors that can align detailed instance descriptions with visual evidence.
@@
Secondly, the agent’s questioning ability remains limited. Reliably disambiguating a target instance from same-category distractors through dialog is still difficult. VLLN-D performs worse on IIGN (20.2%) than VLLN-I on IGN (22.4%), indicating that dialog offers less guidance than a full instruction. We also find that a gap persists between agent-driven and human-driven proactive interaction. Human-Human reaches 93% SR with only about two questions on average, which suggests that a small number of well-chosen queries is sufficient to complete the task. Achieving such behavior requires the agent to exploit its observation history to select maximally informative questions that either guide exploration or shrink the candidate set, placing stronger demands on grounding, planning, and reasoning.
@@ IV-C Additional Insights
Reliability of evaluation and Oracle The benchmark provides a reliable testbed for IIGN, as humans achieve high success rates. The oracle is also dependable, with the Human–Oracle setting attaining performance comparable to Human–Human (91% vs. 93%), though at the cost of more dialog turns (9.72 vs. 2.04 on average). This discrepancy arises for two reasons: some queries fall outside the oracle’s knowledge or response schema, and users interacting with AI systems often seek additional confirmation (e.g., asking a disambiguation question even after the target has been identified). Although not flawless, the oracle is sufficient for evaluating IIGN, as performance in the Agent–Human and Agent–Oracle settings is similar (16% vs. 17%).
@@ V Conclusion
This paper investigates the Interactive Instance Goal Navigation (IIGN) task and introduces the VL-LN benchmark, which includes a long-horizon dataset comprising $\sim$ 41k automatically collected dialogue-augmented trajectories for training, along with an evaluation protocol involving an oracle for agent assessment. This benchmark enables agents to explore long-horizon environments and engage in meaningful dialogs. Our experiments demonstrate that incorporating active dialog significantly improves performance in both the IIGN and IGN tasks, achieving state-of-the-art results. Additionally, we highlight key challenges in the IIGN task and provide some interesting findings from our experiments.